# LLM-ARK: Knowledge Graph Reasoning Using Large Language Models via Deep Reinforcement Learning

### 18 Dec 2023 v1

@@ Abstract
With the evolution of pre-training methods, large language models (LLMs) have exhibited exemplary reasoning capabilities via prompt engineering. However, the absence of Knowledge Graph (KG) environment awareness and the challenge of engineering viable optimization mechanisms for intermediary reasoning processes, constrict the performance of LLMs on KG reasoning tasks compared to smaller models. We introduce LLM-ARK, a LLM grounded KG reasoning agent designed to deliver precise and adaptable predictions on KG paths. LLM-ARK utilizes Full Textual Environment (FTE) prompts to assimilate state information for each step-sized intelligence. Leveraging LLMs to richly encode and represent various types of inputs and integrate the knowledge graph further with path environment data, before making the final decision. Reframing the Knowledge Graph (KG) multi-hop inference problem as a sequential decision-making issue, we optimize our model using the Proximal Policy Optimization (PPO) online policy gradient reinforcement learning algorithm which allows the model to learn from a vast array of reward signals across diverse tasks and environments. We evaluate state-of-the-art LLM(GPT-4) and our method which using open-source models of varying sizes on OpenDialKG dataset. Our experiment shows that LLaMA7B-ARK provides excellent results with a performance rate of 48.75% for the target@1 evaluation metric, far exceeding the current state-of-the-art model by 17.64 percentage points. Meanwhile, GPT-4 accomplished a score of only 14.91%, further highlighting the efficacy and complexity of our methodology. Our code is available on GitHub [^1] https://github.com/Aipura/LLM-ARK for further access.
Keywords KG  $\cdot$ LLM  $\cdot$ DRL
@@ 1 Introduction
With significant progress in large-scale language models (LLMs), researchers have recognized their superior capabilities in the field of natural language processing (NLP)(Liu et al., 2023; Shakarian et al., 2023; Lai et al., 2023). Reasoning ability stands as the most demonstrative of AI intelligence. Recently, to boost the performance of LLMs in reasoning tasks, we noted various optimization strategies adopted by researchers such as Chain of Thought (COT)(Wei et al., 2023) and decomposing subtasks(Kazemi et al., 2023). At the present time, current LLMs reasoning methodologies have seen limited studies in the KG reasoning task. This research endeavor is aimed at addressing this shortfall in the area and contributing to the large-model KG reasoning task.
Knowledge Graphs composed of vertices or entities connected by edges or relations, gaining popularity in knowledge-based systems for its structured disposition. Previous work(Moon et al., 2019; Zhang et al., 2020; Ni et al., 2022; Tuan et al., 2022) mainly relied on supervised learning methods early on. To assess the capabilities of current state-of-the-art (SOTA) LLM: GPT4(OpenAI, 2023), we initially examine the proficiency of LLMs on KG reasoning, as illustrated in Figure 1, determining their potential application in the KG domain. Empirical studies reveal that, despite demonstrating reasonable performance on KG tasks, indicative of their proficiency in managing complex problems, understanding contextual relationships, and utilizing pre-training knowledge, LLMs still present issues and fall short when compared with state-of-the-art models.
@@
There are two main challenges applying LLM-based agents. On the one hand, LLMs suffer from a limited perception of variable reasoning environments. The alignment between LLMs’ knowledge and the environment can be wrong and limit functional competence due to lack of grounding(Carta et al., 2023). If properly grounded, the model’s structure would be both simplified and effective. For KG reasoning tasks, as shown in Figure 1, the agent achieved better scores when provided with as much information about the environment as possible, such as dialog history, inference path history, and all exit paths. Although LLMs are not designed to handle tool use or take actions, Peng et al. (2023); Carta et al. (2023); Du et al. (2023) found that it can be achieved good results in downstream decision making by simply feeding full textual representations as inputs to LLMs.
On the other hand, Yao et al. (2023) indicate that there is a lack of systematic methodologies for consistent model refinement. In other word, LLMs lack the necessary mechanisms for optimizing intermediate reasoning processes in multi-hop reasoning tasks. Manual prompt tuning is still widely used in a lot of the application scenarios. LLM itself is stateless. It has been observed that LLM based agents are easy to run into infinite loops if the states are not handled nicely. Inevitably this methodology runs into the prompt length issues. As the trajectory grows longer, the prompt runs out of spaces. In addition, some restrictions are explicitly specified in the prompt, it is less realistic in knowledge graph environments where a single entity may have more than 100 exit edges. As for supervised fine-tuning, most of the well-performing LLMs are too large to be fit in just one or two GPUs. Most of the issues are caused by the fact that LLMs are not pre-trained or designed with an action-agent application in mind.
Finally, we introduce LLM-ARK, a simple yet effective framework that leverage Large Language Model as Agent to perform Reasoning on Knowledge Graphs, as depicted in Figure 2. We employ LLMs as agent and express the Large model KG inference task as a reinforcement learning sequential decision-making problem, and using a Full-Textual-Environment prompt to aggregate multiscale inputs, we employ LLMs to obtain a richly informative representation of state embeddings. Moreover, our agent architecture does not necessitate access to LLM parameters or gradient propagation through it. Instead, we adopt a policy gradient approach where the Actor LLM functions as part of the environment. This setup allows the model to learn from a vast array of reward signals across diverse tasks and environments. In summary, our contributions are as follows:
  • We assess the capabilities of state-of-the-art LLM: GPT4, on large-scale KG inference datasets and analyze the experimental results in detail to understand the causes of their inferior performance.
@@
  • Furthermore, we update only the parameters of the actors that are part of the LLM agent using the policy gradient method, without accessing the LLM parameters or propagating gradients through them. This approach enables learning from diverse reward signals during interactions with the environment.
@@ 2 Related Work
2.1 KG Reasoning on Dialog Systems && LLMs
@@
KG, due to its structured nature, is an external information source gaining increased popularity in knowledge-based systems. Jung et al. (2020) put forth a model for navigating the conditional graph of a conversation within a KG dialogue system. Leveraging a bidirectional attention flow, the model maximizes the use of structured information from the KG and flexibly determines the extension range of nodes and edges. Zhang et al. (2020) employed a similar approach, using graph attention to navigate the conceptual space of a commonsense KG. Other models, such as Xu et al. (2020), utilize the KG as an external source for controlling coarse-level discourse generation. These innovative uses of KGs facilitate meaningful dialogue generation and nurture topic-focused conversations supported by common sense knowledge. Moon et al. (2019) developed a retrieval system designed to generate responses based on a graph reasoning task. They employed a graph walker to navigate the graph, propelled by the symbol transformation conditions of the dialog context. Expanding on this, Ni et al. (2022) introduced a hierarchical reinforcement learning KG inference model that aggregates multiple inputs utilizing an attention mechanism. This approach instructs the model to reason in one step and then fine-tunes it using a goal-directed approach. Tuan et al. (2022) employed a similar approach, but with a single transformer model that walks directly over large-scale KGs, reasoning over fine-tunable KGs to generate responses. Similarly, Luo et al. (2023) initially create relational paths derived from KGs as high-confidence plans, which are later utilized to extract valid reasoning paths from KGs for confident reasoning.
2.2 LLMs with Deep Reinforcement Learning
@@
The framework carries out exploration and reasoning by pinpointing entities pertinent to a given problem and extracting relevant triples from external knowledge bases. Expanding the scope of the research, Carta et al. (2023) examined large models’ interaction with physical environments, using text to describe these environments, and enhanced them using online reinforcement learning. Yao et al. (2023) employed a special RLHF technique to tailor the model to human preferences, generating beneficial, non-toxic, and safe data for training while also training reward models to evaluate LLMs. Retroformer, a significant improvement over Chain of Thought (COT), is primarily applied to reasoning tasks and uses a unique RLHF methodology. Huang et al. (2022) consistently integrate feedback from diverse sources into the planning language cues of the LLM, thereby enabling it to reason and replan to solve complex problems in both simulated and real-world environments. Zhu et al. (2023) developed robots for open-world environments with generalized capabilities through a large language model grounded in textual knowledge and memory. The LLM utilizes knowledge and memory-based strategies to execute a multilevel disassembly by first breaking down the goal into subgoals via the LLM Decomposer. Then, it breaks down the subgoals into structured actions using the LLM Planner, and finally decomposes the structured actions into a series of actual actions using the LLM Interface. Singh et al. (2022) incorporate Large Language Models to propose a procedural LLM hint structure that facilitates plan generation functionality in contextual environments, robot capabilities, and tasks. In contrast, Shinn et al. (2023) introduce a novel framework, Reflexion, that strengthens linguistic agents through linguistic feedback, rather than updating weights. The Reflexion agent verbally reflects on task feedback signals, and then stores its reflection text in an episodic memory buffer to make better decisions in subsequent trials.
@@ 3.1 Overview
As shown in figure 2, our model has the following main components, FTE(Full-Textual-Environment), LLM(Large Language Model) and RL(Reinforcement Learning). FTE can be seen as state manager, using a Full-Textual-Environment prompt to aggregate multi-scale inputs, updating and maintaining state transfers between itself and the environment, employs a text dictionary format. At first LLMs obtain a richly informative representation of state embeddings. To capture the path embedding information of the KG, we pre-train the KG on TransE(Fan et al., 2014). Instead of feeding in the probability distribution of the action space, our Actor introduces the probability distribution that syncs with the path embedding dimension. Rather than directly introducing the probability distribution of the action space, our Actor feeds the probability distribution along with the path embedding, subsequently eliminating invalid paths (we utilise ’Pad’ for this adaptation process) before outputting a precise and legitimate action. We formulate the large model KG inference task within an online reinforcement learning framework and continuously optimize the decision network based on the collected experience. Finally, we refine the adapter using the Proximal Policy Optimization (PPO) reinforcement learning method, which is based on the obtained reward information.
3.2 Prompt Engineering
As illustrated in Figure 1, we detail the prompting schemes, encompassing the normal prompt and out path aware prompt. To guide LLMs in performing specific dialogue tasks, we can formulate the normal prompt scheme as:
@@ 3.3 LLM-ARK
  • State. $S_{t}\in\mathcal{S}$ is an infinite set of environment states, which encode information stored in Working Memory, including task background $tb$, user query $q$, dialog history $h$, current entity $v_{c}$, path history $ph$, current step $k$, The normal state is represented using a six-tuple: S = (tb, q, h, v_c, ph, k).
  • Observation. The complete state of the environment can be observed. Formally, the observation function $\mathcal{O}=S$.
  • Action. The set of possible actions $\mathcal{A}_{Sm}$ from a state $S_{m}$ consists of all outgoing edges of the vertex $v_{p}$ in G. Formally $\mathcal{A}_{S}=\{e\in E\colon S\}\cup\{(s,\varnothing,s)\}$. This means that each state’s agent can choose one of all output edges $e$ which having the knowledge of the label of the edge $r_{t},r_{t}\in\mathcal{R}$ and destination vertex $v_{t}$.
  • Transition. Depending on the edge selected by the agent at time step $t$, the environment is changed deterministically by updating the state to the new vertex. For single turn dialogue, we update current entity, path history and step. Formally, the transition function $\colon\mathcal{\delta\colon S\times A\rightarrow S}$ is defined by $\mathcal{\delta}(S,A)=(tb,q,h,v\_c^{{}^{\prime}},ph^{{}^{\prime}},k^{{}^{\prime}})$, For multi-turn dialogue wo also need to update user query and dialogue history. $\mathcal{\delta}(S,A)=(tb,q^{{}^{\prime}},h^{{}^{\prime}},v\_c^{{}^{\prime}},ph^{{}^{\prime}},k^{{}^{\prime}})$, where $S=(tb,q,h,c,ph,k)$ and $A=(v_{c},e_{t}))$.
  • Reward. We have a final reward of +1 if the current entity is the target entity $v_{g}$ and -1 otherwise. if $S_{T}=(tb,q^{{}^{\prime}},h^{{}^{\prime}},v\_c^{{}^{\prime}},ph^{{}^{\prime}},k^{{}^{\prime}})$ is the final state, then we have a final reward of +1 if $v\_c^{{}^{\prime}}=v_{g}$ , else -1, i.e., $R(S_{T})=I_{f}\{v_{p}=v_{g}\}$.
@@ 3.3.1 Full Textual Environment
This module (FTE) tracks the agent’s state that captures all essential information in the conversation so far. The same as Prompt Engineering normal prompt.
@@ 3.3.2 Agent
Reference to previous work Carta et al. (2023), we use standard RL practices by adding action heads - a Multi-Layer Perceptron (MLP) with —A— outputs - on top of the LLM. Thus, we can use only pretrained operations from the LLM and leverage language modeling heads’ prior, this method is robust to any action space and can thus be used on any textual environment with no change. We design a randomized nonstationary policy $\pi=(d_{0},d_{1},...,d_{t-1})$, where $d_{t}=P(\mathcal{A}_{S_{t}})$ is a policy at time step t. Agent has two components, one of which is LLM and the other is PA-MLP.
LLM We initially utilize a LLM to encode the state $S_{t}=(S_{t-1},A_{t-1},V_{t})$ into a continuous vector $s_{t}\in\mathcal{\mathbb{R}}^{2d}$. As a rule of thumb, for BERT models the cls token is used to represent the semantics of the whole sequence, while standard transformers and GPT-like LLMs use the embedding of the last token. We used the model on huggingface hub as well as the code to get the sequence vector representation [^2] https://huggingface.co. It is defined by:
@@
PA-MLP
Instead of just adding an MLP with a single output for the value on top of the last layer of the first decoder block as in the conventional multicategorization task, to enhance the ability of the large model to perceive the environment, we further fused the hidden state after the MLP with the Knowledge Graph exit path information, called PA-MLP(Path Aware MLP). Recall that each possible action represents an outgoing edge $e$ with information about the edge relation label $r_{l}$ and the target vertex/entity $v_{d}$. So the embedding for each $A\in A_{St}$ is $[r_{l};v_{d}]$, and stacking the embeddings for all outgoing edges we get the matrix $A_{t}$. The network taking this as input is parameterized as a three-layer feed-forward network (MLP) with $\tanh$ nonlinearity, which takes the FTE representation $FTE_{t}$ and the embedding for the outgoing paths embedding $e_{g}$ and outputs a probability distribution over the possible actions from which a discrete action is sampled. Note that the dimension of the MLP output hidden state must be equal to the dimension of the path embedding. Finally, formulated as:
$\mathbf{h}_{\mathbf{t}}$ $=\mathbf{A}_{\mathbf{t}}\left(\mathbf{W}_{\mathbf{3}}\operatorname{tanh}\left(\mathbf{W}_{\mathbf{2}}\operatorname{tanh}\left(\mathbf{W}_{\mathbf{1}}\left(s_{t}\right)\right)\right)\right)$ (4)

### 04 Feb 2024 v2

# Evaluating and Enhancing Large Language Models for Conversational Reasoning on Knowledge Graphs

@@ Abstract
The development of large language models (LLMs) has been catalyzed by advancements in pre-training techniques. These models have demonstrated robust reasoning capabilities through manually designed prompts. In this work, we evaluate the conversational reasoning capabilities of the current state-of-the-art LLM (GPT-4) on knowledge graphs (KGs). However, the performance of LLMs is constrained due to a lack of KG environment awareness and the difficulties in developing effective optimization mechanisms for intermediary reasoning stages. We further introduce LLM-ARK, a LLM grounded KG reasoning agent designed to deliver precise and adaptable predictions on KG paths. LLM-ARK leverages Full Textual Environment (FTE) prompt to assimilate state information within each reasoning step. We reframe the challenge of multi-hop reasoning on the KG as a sequential decision-making task. Utilizing the Proximal Policy Optimization (PPO) online policy gradient reinforcement learning algorithm, our model is optimized to learn from rich reward signals. Additionally, we conduct an evaluation of our model and GPT-4 on the OpenDialKG dataset. The experimental results reveal that LLaMA-2-7B-ARK outperforms the current state-of-the-art model by 5.28 percentage points, with a performance rate of 36.39% on the target@1 evaluation metric. Meanwhile, GPT-4 scored 14.91%, further demonstrating the effectiveness of our method.
@@ 1 Introduction
With significant progress in large language models (LLMs), researchers have recognized their superior capabilities in the field of natural language processing (NLP)(Liu et al., 2023; Shakarian et al., 2023; Lai et al., 2023). Reasoning ability stands as the most demonstrative of AI intelligence. Recently, to boost the performance of LLMs in reasoning tasks, we noted various optimization strategies adopted by researchers such as Chain of Thought (COT)(Wei et al., 2023) and decomposing subtasks(Kazemi et al., 2023). Currently, the reasoning method of LLMs have received limited attention in the conversational KG reasoning task. This research aims to address this gap in the field.
Knowledge Graphs composed of vertices or entities connected by edges or relations, gaining popularity in knowledge-based dialogue systems for its structured disposition. Conversational reasoning models are able to traverse the KG based on conversational context to introduce diverse entities and attributes to make replies more engaging, as well as to improve the logic of the model to mitigate illusions(Rawte et al., 2023; Dong et al., 2022). Previous work(Moon et al., 2019; Zhang et al., 2020; Ni et al., 2022; Tuan et al., 2022) mainly relied on supervised learning methods. To assess the capabilities of current SOTA LLM: GPT-4(OpenAI, 2023), we initially examine the proficiency of LLMs on KG reasoning, as illustrated in Figure 1, determining their potential application in the KG domain. Empirical studies reveal that, despite demonstrating reasonable performance on KG tasks, indicative of their proficiency in managing complex problems, understanding contextual relationships, and utilizing pre-training knowledge, LLMs still present issues and fall short when compared with state-of-the-art models.
@@
There are two main challenges applying LLM-based agents. On the one hand, LLMs suffer from a limited perception of variable reasoning environments. The alignment between LLMs’ knowledge and the environment can be wrong and limit functional competence due to lack of grounding(Carta et al., 2023). If properly grounded, the model’s structure would be both simplified and effective. For KG reasoning tasks, as shown in Figure 1, the agent achieved better scores when provided with as much information about the environment as possible, such as dialog history, inference path history, and all exit paths. Although LLMs are not designed to take actions, Peng et al. (2023); Carta et al. (2023); Du et al. (2023) found that it can be achieved good results in downstream decision making by simply feeding full textual representations as inputs to LLMs.
On the other hand, Yao et al. (2023) indicate that there is a lack of systematic methods for consistent model refinement. In essence, LLMs fall short in possessing essential mechanisms for optimizing intermediary reasoning processes in multi-hop reasoning tasks. This is mainly attributed to the fact that manual prompt tuning is widely used in many application scenarios. It has been observed that LLM-based agents can easily fall into infinite loops if state is not handled properly, and inevitably run into prompt length problems when the trajectory becomes longer. In addition, the design of prompt is also a challenge because an entity may have more than 100 exit edges, all of which are formatted into prompt which is impractical in a knowledge graph environment. LLMs often encounter these issues because they are not designed or trained for action-agent applications.
We introduce LLM-ARK, an effective framework that employs LLM as an Agent for Reasoning on Knowledge Graphs. We employ LLMs as agent and express the Large model KG inference task as a reinforcement learning sequential decision-making problem, and using a Full-Textual-Environment prompt to aggregate multiscale inputs. Moreover, our agent architecture does not necessitate access to LLM parameters or gradient propagation through it. Instead, we adopt a policy gradient approach where the Actor LLM functions as part of the environment. This configuration enables the model to learn from diverse reward signals across varied tasks and environments. In summary, our contributions are as follows:
  • We assess the capabilities of state-of-the-art LLM: GPT-4, on large-scale KG inference datasets and analyze the experimental results in detail to understand the causes of their inferior performance.
@@
  • Furthermore, we update only the parameters of the PA-MLP that are part of the our agent using the policy gradient method, freezing the parameters of the LLM. This approach enables learning from diverse reward signals during interactions with the environment and improves the efficiency of training.
@@ 2 Related Work
2.1 KG Reasoning on Dialog Systems
@@
Given its structured nature, Knowledge Graphs are becoming an increasingly popular external information source in knowledge-based systems. Moon et al. (2019) developed a retrieval system designed to generate responses based on a graph reasoning task. They employed a graph walker to navigate the graph, propelled by the symbol transformation conditions of the dialog context. Jung et al. (2020) utilizing graph attention techniques to navigate the conditional graph of a conversation within a KG dialogue system. The model computes an incoming attention fow to represent entities and an outgoing attention fow to select KG paths. However, this approach cannot be extended to long KG path prediction due to the exponential increase in computational complexity. Ni et al. (2022) introduced a hierarchical reinforcement learning KG inference model that aggregates multiple inputs utilizing an attention mechanism. This approach instructs the model to reason in one step and then fine-tunes it using a goal-directed reinforcement learning. Tuan et al. (2022) employed a single transformer model that walks directly over large-scale KGs, reasoning over fine-tunable KGs to generate responses. Similarly, Luo et al. (2023) initially create relational paths derived from KGs as high-confidence plans, which are later utilized to extract valid reasoning paths from KGs for confident reasoning. Sun et al. (2023) leverage KGs to augment LLMs for deep and responsible reasoning. The framework explores and infers by identifying entities relevant to a given question and retrieving relevant triples from external KGs. This iterative process generates multiple inference paths until enough information is gathered to answer the question or maximum depth is reached.
2.2 LLMs with Reinforcement Learning
@@
Reinforcement learning and large models are divided into two main aspects of the combination, the first aspect further improves the ability of LLM to understand and follow user instructions through reinforcement learning based methods(Ouyang et al., 2022). Yao et al. (2023) employed a special RLHF technique to tailor the model to human preferences, generating beneficial, non-toxic, and safe data for training while also training reward models to evaluate LLMs. Retroformer, a significant improvement over Chain of Thought (COT), is primarily applied to reasoning tasks and uses a unique RLHF method. Shinn et al. (2023) introduce a novel framework, Reflexion, that strengthens linguistic agents through linguistic feedback, rather than updating weights. The Reflexion agent verbally reflects on task feedback signals, and then stores its reflection text in an episodic memory buffer to make better decisions in subsequent trials.
The second aspect is to further improve the applicability of LLM on real-world tasks through a reinforcement learning-based approach, since training a LLM public NLP task/dataset can only cover a small portion of the real world, and reinforcement learning can train LLM-based intelligences to explore the realization of various real-world goals. Carta et al. (2023) studied LLMs interaction with physical environments. Using an interactive textual environment designed to study a series of spatial and navigational tasks and using online reinforcement learning to improve its performance to solve goals. Huang et al. (2022) consistently integrate feedback from diverse sources into the planning language cues of the LLM, thereby enabling it to reason and replan to solve complex problems in both simulated and real-world environments. Singh et al. (2022) propose a procedural LLM hint structure that facilitates plan generation functionality in contextual environments, robot capabilities, and tasks.
@@ 3.1 Overview
As shown in Figure 2, our model has the following main components, FTE (Full-Textual-Environment), LLM (Large Language Model) and RL(Reinforcement Learning). FTE can be seen as state manager, using a Full-Textual-Environment prompt to aggregate multi-scale inputs, updating and maintaining state transfers between itself and the environment. At first LLMs obtain a richly informative representation of state embeddings. To capture the path embedding information of the KG, we pre-train the KG on TransE(Fan et al., 2014). Rather than directly introducing the probability distribution of the action space, our Actor feeds the probability distribution along with the path embedding, subsequently eliminating invalid paths (we utilise ’Pad’ for this adaptation process) before outputting a precise and legitimate action. We formulate the large model KG inference task within an online reinforcement learning framework and continuously optimize the decision network based on the collected experience in replay buffer. Finally, we refine the adapter using the Proximal Policy Optimization (PPO) online reinforcement learning method. In this section, we will first describe the method used to evaluate GPT-4, and then present each of the modules of our model in turn.
3.2 Manual Prompt Tuning
As illustrated in Figure 1, we detail the prompting schemes, encompassing the standard prompt, normal prompt and out path aware prompt. To guide LLMs in performing specific dialogue tasks, we can formulate the standard prompt and normal prompt scheme as:
@@ 3.3 LLM-ARK
  • State. $S$ is an infinite set of environment states, which encode information stored in Working Memory, including task background $tb$, user query $q$, dialog history $h$, current entity $v_{c}$, path history $ph$, current step $t$, The normal state is represented using a six-tuple: $S=(tb,q,h,v_{c},ph,t)$.
  • Observation. The complete state of the environment can be observed. Formally, the observation function $O=S$.
  • Action. The set of possible actions $\mathcal{A}$ from a state $S$ consists of all the environment information. Formally $\mathcal{A}_{s}=\{e\in E\colon S\}\cup\{(s,\varnothing,s)\}$. This means that each state’s agent can choose one of all output edges of the current entity.
  • Transition. Depending on the edge selected by the agent at time step $t$, the environment is changed deterministically by updating the state to the new environment. For single turn dialogue, we update current entity, path history and step. Formally, the transition function $\colon\mathcal{\delta\colon S\times A\rightarrow S}$ is defined by $\mathcal{\delta}(S,A)=(tb,q,h,v_{c}^{{}^{\prime}},ph^{{}^{\prime}},t^{{}^{\prime}})$, For multi-turn dialogue wo also need to update user query and dialogue history. $\mathcal{\delta}(S,A)=(tb,q^{{}^{\prime}},h^{{}^{\prime}},v_{c}^{{}^{\prime}},ph^{{}^{\prime}},t^{{}^{\prime}})$, where $S=(tb,q,h,v_{c},ph,t)$.
  • Reward. We have a final reward of +1 if the current entity is the target entity $v_{g}$ and -1 otherwise. if $S_{t}=(tb,q^{{}^{\prime}},h^{{}^{\prime}},v_{c}^{{}^{\prime}},ph^{{}^{\prime}},k^{{}^{\prime}})$ is the final state, then we have a final reward of +1 if $v_{c}^{{}^{\prime}}=v_{g}$ , else -1.
@@ 3.3.1 Full Textual Environment
This module tracks the agent’s state that captures all essential information in the conversation so far. FTE is a text dictionary structure, the same as Prompt Engineering’s normal prompt format.
@@ 3.3.2 Agent
Inference to previous work Carta et al. (2023), we use standard RL practices by adding action heads - a Multi-Layer Perceptron (MLP) on top of the LLM. Thus, we can use only pretrained operations from the LLM and leverage language modeling heads’ prior, this method is robust to any action space and can thus be used on any textual environment with no change. Agent has two components: LLM and PA-MLP.
LLM We initially utilize a LLM to encode the state $S$ into a continuous vector $s\in\mathcal{\mathbb{R}}^{2d}$. As a rule of thumb, for BERT models the cls token is used to represent the semantics of the whole sequence, while standard transformers and GPT-like LLMs use the embedding of the last token. We used the model on huggingface hub as well as the code to get the sequence vector representation [^1] https://huggingface.co. It is defined by:
@@
PA-MLP Instead of just adding an MLP with a single output for the value on top of the last layer of the first decoder block as in the conventional multicategorization task, to enhance the ability of the large model to perceive the environment, we further fused the hidden state after the MLP with the Knowledge Graph exit path information, called PA-MLP (Path Aware MLP). Recall that each possible action represents an outgoing edge $e$ with information about the edge relation label $r_{l}$ and the target vertex/entity $v_{d}$. So the embedding for each $A_{t}$ is $[r_{l};v_{d}]$, and stacking the embeddings for all outgoing edges we get the matrix $A_{t}$. The network taking this as input is parameterized as a three-layer feed-forward network (MLP) with $\tanh$ nonlinearity, which takes the FTE representation $s$ and the embedding for the outgoing paths embedding and outputs a probability distribution over the possible actions from which a discrete action is sampled. The dimension of the MLP output hidden state is equal to the dimension of the path embedding. Finally, formulated as:
$\mathbf{h}_{\mathbf{t}}$ $=\mathbf{A}_{\mathbf{t}}\left(\mathbf{W}_{\mathbf{3}}\left(\operatorname{tanh}\left(\mathbf{W}_{\mathbf{2}}\left(\operatorname{tanh}\left(\mathbf{W}_{\mathbf{1}}\left(s_{t}\right)\right)\right)\right)\right)\right)$ (4)